HVAMS03: Logbook Entries

HVAMS03: Site 2 Messages, 17 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
144 Wed 22-Oct-2003st2: sdr installed, eve problem2militzer
143 Wed 22-Oct-2003st2: FastData Outage 2militzer
140 Tue 21-Oct-2003st2: kh2o cleaned, routine2militzer
122 Wed 15-Oct-2003st2: CO2 Noise, Cal-table, Antenna moved2militzer
118 Mon 13-Oct-2003St2: Swap SOAD 'plate' (had noise during 'goes')2delany
114 Sun 12-Oct-2003St2: CO2 cal, routine2militzer
101 Mon 06-Oct-2003Site 2 visit, Oct 62holden
90 Fri 03-Oct-2003Site 2 visit, Oct 22holden
88 Thu 02-Oct-2003Site 2 Q soil service, Oct 12holden
85 Wed 01-Oct-2003Site 2 visit, Sept 302holden
75 Sat 27-Sep-2003Site 2 visit, Sept 272oncley
70 Thu 25-Sep-2003Site 2 visit, Sept 252horst
57 Mon 22-Sep-2003Site 2 visit, Sept 222horst
56 Mon 22-Sep-2003Site 2 visit, Sept 212horst
40 Mon 15-Sep-2003Site 2 visit, Sept 142horst
30 Fri 12-Sep-2003Visit: PSPin fixed...2militzer
16 Mon 11-Aug-2003Site 2 CR10X calibration file2semmer

144: EVE, Site 2, Wed 22-Oct-2003 08:58:38 EDT, st2: sdr installed, eve problem
10/21/03 14:40-15:00

SDR Data: Inserted new CF and verified that it was working OK.

EVE problem: on 10/21 at 0Z the RAMDISK DIR. STRUCT. FAILED!
	This was shown evidenced on syslog.log and on console
	screen in 'failure to access'.
	At that point any 'dir' or 'dir /card' failed,
	DAQ fortunately was running, including sdr output (when good media!).
	'REBOOT' RECOVERED full ops and the directory structure was
	actually OK.

	Second/third time this has been seen: systems getting older
	which is a problem, however:
1) EVE code must be examined closely for this problem (ie interrupt
	contention on disk access, etc.) and auto recovery procedures.
2) DCOM332 boards must be fully exercized and tested, possibly checked
	by manuf. as well.   Using 'in-house' software, temp, etc.
143: Status, Site 2, Wed 22-Oct-2003 08:46:29 EDT, st2: FastData Outage
SDR Fast Data Outage From 10/14 to 10/20:

The upload of the 'sdr fastdata' compact flash recovered on
10/20 failed.  After several checks, including back on a 'sdr'
we found:    The MEDIA Itself FAILED !!!

Sandisk 06/03-3 had been used extensively during pre-proj tests
however this indicates 2 things:
1) During project:	While servicing and esp. when inserting
   a blank media, make sure data is being written successfully
   by watching file sizes grow.  (problem appeared in dir part.)
   I didn't do that on st2 on 10/14 and the Tern dutifully created
   file and happily chugged away even though the CF itself wasn't
2) After project:	Perform CF life cycle testing to better
   anticipate replacement schedule.
140: Site_visit, Site 2, Tue 21-Oct-2003 10:26:00 EDT, st2: kh2o cleaned, routine
10/20/03 12:20-12:55

kh2o cleaned: everything was wet from dew, output went from .3mv to ~660mV

Soil sample taken

RG: ok but the bucket was essentially full from the last rain event,
	which means it'll tip early next time.

SDR/metppf uploaded:  NOTE: it appears the sdr data files were corrupt
	and this needs to be checked on-site. Data loss from 10/15-20.

RMY BD still set =000
122: Site_visit, Site 2, Wed 15-Oct-2003 09:35:31 EDT, st2: CO2 Noise, Cal-table, Antenna moved
10/14/03 15:40-16:30, and 16:55-17:25

CO2/O3,etc Noise Problem: (see also entry 118)
The occurance of 5 min noise spikes in the AD data for this station indicated
 GOES RF interference.
The Yagi antenna was relocated from the SE leg of the PAM tripod to the W leg
 and then aligned to the original azimuth and tilt pointing.

CO2 Cal Tables:
The CO2 sensor data had been showing steps as the calibration switched between
different calibratin tables when the temperature hovers around 10-degC.
SteveO thought it would be useful to change the temp. threshold value to
keep the unit from switching tables:
>st	stop unit
>tr0 29500 3080 0 0 0.7479	original values
>tr1 23000 2830 1 1 0.9035
>tr2 15300 2630 2 2 1.0590
Per SO changed tr0 from 3080 (ie 308degK) to 3280 and restarted.
Later returned per SO to change tr1 which he thought more appropriate.
That was put to 3030 (ie 303degK) however when doing that the unit would
not 'go' again, apparently due to inability to hit this temp. on the chiller.
Final Result:	Returned tr0 and tr1 to original values so it'll work!

118: Site_visit, Site 2, Mon 13-Oct-2003 17:16:12 EDT, St2: Swap SOAD 'plate' (had noise during 'goes')
10/13/03 15:20-16:10 local

Purpose was to investigate Data glitch on AD during Xmit.
Suspected that the SOAD had the 'single-ended' jumper still on the
board but that was not the case.

'Serializer/AD' plates swapped 5 times between 15:20 and 15:30, while watching
data on PC.

At 15:45 spare replacement plate installed and fast data accumulated
for 20 minutes before being downloaded for exam back at base.
NOTE: at one point we observed microbaro voltage reported at 180mV but with
      spare plate installed it was reporting at 1180mV....This could have
      been due to the ad cable being plugged in slightly incorrectly when
      we were using the original 'plate'.   
NOTE: We did observe the 1180mV jump up by 5mV during xmit, however, this
      is far less than what was seen in the fastout data (I think).....
	wrong!...a check of the 'caled' data back at the base shows
	that the step is still there.  It equals about .1mB but is
	very clear in the data.
NOTE: The CO2 pump audibly changes during xmit, indicating reduced airflow
      perhaps and that could account for the increased 5mV signal ....maybe?
	We should check where the pump is plugged in and
	maybe try an 'independent' battery for it.
NOTE: The noise is also on O3, Vbatt, Tbatt but not Tsfc.
	This implies location adjacent to the antenna is the culprit.
	St2 has a yaggi located on the SE leg, immediately in line with
	the microbaro, O3, etc.   Cable shielding is the other possibility
	and I don't know if SUNY put the shield on or not.

Datascope used to check boom angles
    RMY boom = 359 degrees,    Sonic boom = 91 degrees

114: CO2, Site 2, Sun 12-Oct-2003 16:16:34 EDT, St2: CO2 cal, routine
Arrived 15:06, leave 15:52

CO2 cal performed by Sachai
Radiometers cleaned
Soil sample taken
Trime measured 

SDR data uploaded
Metppf data uploaded

101: Site_visit, Site 2, Mon 06-Oct-2003 16:51:18 EDT, Site 2 visit, Oct 6
Site 2 visit, Oct 6, SPO, KDH, & Sakai
Arrived 1440 EDT

1440 shadow test, 2 cycles
CO2 reading approx. 800ppm, lights on RMT are green
Load edited S02.0160.dat with new scale factors.
Cleaned rads.
1450 TRIME: 43.0, 44.1, 43.1%
core taken
1453 grabbed SDR

90: Site_visit, Site 2, Fri 03-Oct-2003 09:54:02 EDT, Site 2 visit, Oct 2
Site 2 visit, Oct 2, SPO, KDH, & Sakai
Arrived 1210 EDT

Arrived format was 160. Used that format.

Installed uBaro, Ozone, & CO2.  
1305: Moved inlets of Ozone & CO2 nearer port.

Grabbed SDR to check user sensors.

1315: left site

88: Site_visit, Site 2, Thu 02-Oct-2003 09:43:04 EDT, Site 2 Q soil service, Oct 1
Site 2 visit, Oct 1, SPO & KDH
Arrived 1410 EDT

Visit site to fix Q soil.
Found cut cable.  Spliced with soldering iron.  Problem fixed.

1455 left site

85: Site_visit, Site 2, Wed 01-Oct-2003 11:21:18 EDT, Site 2 visit, Sept 30
Site 2 visit, Sept 30, SPO & KDH
Arrived 1514 EDT

Cleaned krypton 140mV->400
Mast down 1519->1524

Grabbed SDR
Copied metppf

1535: Trime 38.9; 36.9; 36.5%
1535: soil core sample taken

1529: Cleaned rad, no paddle test due to cloudy skies

1514-1612: Rain gauge calibration started 303-414

Garmin 42deg 18.423' 73deg 50.906' 47m(16')


1612 left site

75: Site_visit, Site 2, Sat 27-Sep-2003 15:18:49 EDT, Site 2 visit, Sept 27
Site 2 visit, Sept 27, SPO
Arrived 1300 EDT; rain now heavy at times, but some clearing.
Walked the path from the road (really muddy, but found a Leatherman!)

Hoped to replace net dome, but one screw head stripped, so had to replace
entire sensor.  Old: 91096;  New: 94196 (had replaced dome since pulled from 
Disconnected old; carried back to van; brought new to site; entered new coef;
mounted new; connected new.  Thus, there shouldn't be any non-zero data with
bad coefs. 

Grabbed SDR.

1353: Trime: 34.9; 38.3; 40.2%

GARMIN: 42deg 18.418'N; 73deg 50.901'W; 40m (acc 17ft)

- didn't clean krypton
- didn't clean rads
- didn't grab metppf (didn't want to use laptop in rain or keep box open)
- didn't take core (would have meant another sloppy trip)

70: Site_visit, Site 2, Thu 25-Sep-2003 17:11:46 EDT, Site 2 visit, Sept 25
Site 2 visit, Sept 25, SPO and Tom
Arrived 1200 EDT, able to drive to site

Downloaded metppf files and swapped fastout media

SPO collected soil sample at 1206
Trime = 40.6%
	36.8% at 1213

Did not lower mast to clean kh2o; did not clean radiometers

57: Radiation, Site 2, Mon 22-Sep-2003 18:00:02 EDT, Site 2 visit, Sept 22
Site 2 visit, September 22, Kurt and Tom

1220 EDT, Kurt attempted to replace Rnet upper dome.   He could not
loosen the screws.  Needs to replace Rnet with spare and replace dome
in lab.

56: Site_visit, Site 2, Mon 22-Sep-2003 17:52:28 EDT, Site 2 visit, Sept 21
Site 2 visit, September 21, Kurt and Tom
Arrived 1735 EDT

kh2o output = 54 mV; s/n 1133
lowered mast at 1737; raised at 1739
cleaned kh2o; output = 246 mV
Copied new KRY1133s.DAT to EVE and /card
Copied new hvamlogger.all to EVE
restarted EVE

Cleaned radiometers at 1748 EDT
Rnet has scratch and perhaps a hole in top dome

Copied local data storage and swapped fastout

Soil sample at 1754 EDT; Qsoil = 60.03%
Trime reading = 32.5%	no pilot tool
                36.2%	no pilot tool
		38.7%	no pilot tool

40: Site_visit, Site 2, Mon 15-Sep-2003 15:33:46 EDT, Site 2 visit, Sept 14
Visit to station 2, Sept 14, Kurt and Tom
Arrived 1235 EDT
Left    1430 EDT

Mast down 1245 EDT
Mast up   1303 EDT

KH2O output = 5 mV; cleaned; output = 54 mV

RMY BoomDir = 000 deg
Baro s/n = NCAR002
Checked all other s/n's; okay with logbook #26

Downloaded SDR fastout and local data storage
Downloaded logger.all config to station

Took soil sample at 1345 EDT; 3-6 cm depth
Tested rain gauge with 10(?) tips; read 11 tips on Panasonic

Kurt installed new CS615 at 5 cm depth; reads 48% on Panasonic

Time check okay

Kurt took site photos, starting and ending with EVE box #2.
He took 8 photos from N to NE to E to SE, etc.

30: Site_visit, Site 2, Fri 12-Sep-2003 16:20:07 EDT, Visit: PSPin fixed...
9/11/03 ~18Z
Photos Taken
PSPin wiring inside logger fixed (loose wire)
CS615 confirmed bad
Upload metppf / sdr data
New Config downloaded.
16: Loggers, Site 2, Mon 11-Aug-2003 14:52:56 MDT, Site 2 CR10X calibration file
Q91096-A+       NET     00      8.57
A-      x       01      10.53
P31976  PygIn   02      260.484501
P31977  PygOut  03      268.096515
K970379 PspIn   04      45.51661356
K940181 Psput   05      82.3723229
H993561 HFT     06      37.2
x       SlDn    07      0.0
Tcin-A0 x       10      2500
A1      x       11      3.354748
A2      x       12      .279406
A3      x       13      3.926898
A4      x       14      202.8886
Tdin-A0 x       15      2500
A1      x       16      3.053435
A2      x       17      .270616
A3      x       18      3.557806
A4      x       19      167.4648
Tcot-A0 x       20      2500
A1      x       21      3.355048
A2      x       22      .279392
A3      x       23      3.969477
A4      x       24      201.0222
Tdot-A0 x       25      2500
A1      x       26      3.052511
A2      x       27      .270654
A3      x       28      3.516783
A4      x       29      176.0205
NCAR002-A0      Tsoil   30      -243.557624
A1      x       31      234.596987
A2      x       32      7.741747
x-A0    LevelX  33      0.0
A1      x       34      0.0
A2      x       35      0.0
A3      x       36      0.0
A4      x       37      0.0
A5      x       38      0.0
x-A0    LevelY  39      0.0
A1      x       40      0.0
A2      x       41      0.0
A3      x       42      0.0
A4      x       43      0.0
A5      x       44      0.0
002-A0  CS615   45      -.187
A1      x       46      .037
A2      x       47      .335
TP01    timeout 50      180
levels  x       98      0
baud    x       99      1